Manuel Kauers Hagenberg, 19. März 2003 RISC-Linz

Word Order in German Sentences

Main Clauses and Subclauses

Main Clauses – Hauptsätze

• Example:

Mann, m – man geben – give Kind, n – child Schlüssel, m – key Mutter, f – mother

Der Mann gibt dem Kind den Schlüssel der Mutter.

• This sentence contains all cases: der Mann is nominative, dem Kind is dative, den Schüssel is accusative, and der Mutter is genitive (the mother’s key).

• There are other admissible arrangements of this sentence, like

Dem Kind gibt der Mann den Schlüssel der Mutter. Den Schlüssel der Mutter gibt dem Kind der Mann. Den Schlüssel der Mutter gibt der Mann dem Kind. …

• They all have the same meaning (modulo emphasis) and might be used interchangable.

• Every thinkable order of the “building blocks” der Mann, gibt, dem Kind, den Schlüssel der Mutter is admissible provided that the verb gibt comes at the second position.

• This is like planets which are moving around the sun (the verb)

• Note that den Schlüssel der Mutter is only one building block, because der Mutter refers to den Schlüssel, not to the verb gibt.

• In a certain sense, this building block is like a planet with a moon, because also der Mutter can “rotate” around den Schlüssel (then hiding the article): You can say der Mutter Schlüssel — this is correct but not common.

• Rule: In every well formed german sentence, the verb comes at the second position. As a rule of thumb, the position of everything else is arbitrary.

• You can also put the verb at the first position, then you get a question. If the verb is at any other position, then it’s a mistake.

• If the verb is split into parts, then the above rules apply to the conjucated part, and the splitted part goes always to the end.

Exercise

• Separate the building blocks of the following examples by vertical bars.

Morgen – tomorrow

stattfinden – take place kein – no Deutschkurs, m – German class Apfel, m – apple fallen – fall nicht – not weit – far vom ∼ from Stamm – bole or clan besonders – especially gut – good

Morgen findet kein Deutschkurs statt. Der Apfel fällt nicht weit vom Stamm.^{1} Frau Mayrs Englisch ist nicht besonders gut.

• Find all possible arrangements, without questions, of the sentence

sprechen – speak

im – in the über – about Arbeit, f – work Institut, n – institute

Franz spricht im Forum über die Arbeit im Institut.

(Hint: I found six.) In general, if there are n building blocks, how many possible arrangements are there? ;-)

• If there are so many right possibilities to form a sentence, how should you decide which one to choose?

Subclauses – Nebensätze

• If A and B are sentences, then so are

A, und B  and
A, oder B  or
Entweder A, oder B  either A or B
A, denn B  since/because
A, but B  aber
Nicht A, sondern B  Not A, but B
Nicht nur A, sondern auch B Not only A, but also B

• Entweder, Nicht, Nicht nur may or may not count as a building block of A and influence its word order accordingly. Entweder usually does, the others usually don’t.

• Factorisation is possible: Instead of

Franz spricht heute und Heinrich spricht morgen.

you can say

Franz spricht heute und Heinrich morgen.

• These are main clause connections. They are boring…

• We consider subclauses that start with subjunctions. They are summarized in the following table:

1German saying expressing that children’s behaviour is usually similar to that of their parents. (Note the double meaning of Stamm!)

weil  because  obwohl  although
wenn  if (cond.)  ob  if, whether
während while  nachdem after
bevor  before  dass/daß that
als  as (temp.)  falls  in case

• In a subclause, the subjunction takes the first position, and the verb must be put at the very end. There is no splitting of verbs.

können – can arabisch – arabic er – he aus – from Ägypten – egypt kommen – come laut – loud hart – hard arbeiten – work gehen – go nach Hause – (to) home fünf Uhr – five o’clock

Mohammed kann arabisch, weil er aus Ägypten kommt. Es ist laut, während sie hart arbeiten. Er sagt, daß morgen kein Deutschkurs stattfindet. Sie geht nach Hause, wenn es fünf Uhr ist.

• Subclauses can never stand alone. There must be at least one main clause per sentence.

• Subclauses can of course be iterated:

sagen – say

am (an dem) – at the Telefon, n – phone sie – she zum (zu dem) – to the krank – ill

Nino sagt am Telefon, daß sie nicht zum Deutschkurs kommt, weil sie krank ist.

• Clauses can be nested. A subclause can be put at the beginning, at the end, or even in the middle of another subclause:

Weil sie krank ist, geht sie nach Hause. Sie geht, weil sie krank ist, nach Hause. Sie geht nach Hause, weil sie krank ist.

Note that a subclause counts as a building block of the surrounding clause, and may influence its word order.

• Also the question words like wer (who) may serve as subjunctions.

du – you

wissen – know sicher – certainly heute – today fehlen – miss

Du weißt sicher, wer heute fehlt.

Such subclauses are called implicit question clauses (indirekte Fragesätze)

• The articles serve as subjunctions when forming relative clauses (Relativsätze). Relative clauses are put directly behind the noun to which they refer. If there is no abiguity, they may also be put at the end of the clause.

Hund, m – dog bellen – bark beißen – bite

Hunde, die bellen, beißen nicht.^{2}

Exercise

• Build from the vocabulary you know an example for a sentence with subclause to each of the subjunctions from the table. (Omit nachdem and bevor; they require tenses.)

2German saying.

• Build from the vocabulary you know a sentence with more than one subclause.

• In English there is no distinction between implicit question clauses and relative clauses, both start with who, which, whose, etc. What is the difference in German? Give examples for both kind of clauses.

• Formulate the following questions indirectly using phrases like Ich weiss (I know), Es ist egal/Es spielt keine Rolle (It doesn’t matter), Ich kann mir denken (I can imagine), etc.

was – what machen – do wohnen – live/stay gehen – go gleich – soon zum – to the Seminar, n – seminar

Was machst Du morgen? Wo wohnst Du? Wer geht gleich zum Unimarkt? Ist morgen Seminar?^{3}

3Hint: use ob